### 06 Feb 2024 v2

# On the action of the symmetric group on the free LAnKe

Abstract.
A LAnKe (also known as a Filippov algebra or a Lie algebra of the $n$-th kind) is a vector space equipped with a skew-symmetric $n$-linear form that satisfies the generalized Jacobi identity. Friedmann, Hanlon, Stanley and Wachs have shown that the symmetric group acts on the multilinear part of the free LAnKe on $12n-$ 1 generators as an irreducible representation. They have claimed that the multilinear component on $23n-$ 2 generators decomposes as a direct sum of two irreducible symmetric group representations. This paper provides a proof of the later statement.
Key words and phrases:
Specht module, symmetric group, LAnKe, n-ary algebra, Filippov algebra
1991 Mathematics Subject Classification:
05E10, 20C30, 20G05
1. Introduction
Throughout this paper we work over a field $\mathbb{K}$ of characteristic zero.
Since the mid 1980’s various $n$-ary generalizations of Lie algebras have been introduced and studied, see the review by de Azcárraga and Izquierdo [3]. Among those are the Filippov algebras [7], which are also called LAnKes [5].
Definition 1.1 ([5, Definition 1.2]).
A Lie algebra of the $n$-th kind (LAnKe or Filippov algebra) is a $\mathbb{K}$-vector space $\mathcal{L}$ equipped with an $n$-linear bracket $[-,-,\dots,-]:\mathcal{L}^{n}\to\mathcal{L}$ such that for all $x_{1},\dots,x_{n},y_{1},\dots,y_{n-1}\in\mathcal{L}$,
  1. (1)
    $[x_{1},x_{2},\dots,x_{n}]=sign(\sigma)[x_{\sigma(1)},x_{\sigma(2)},\dots,x_{\sigma(n)}$] for every $\sigma\in\mathfrak{S}_{n}$, and
  2. (2)
    the following generalized Jacobi identity (GJI) holds
    (1.1) $[[x_{1},x_{2},\dots,x_{n}],y_{1},\dots,y_{n-1}]$
    $=\sum_{i=1}^{n}[x_{1},x_{2},\dots,x_{i-1},[x_{i},y_{1},\dots,y_{n-1}],x_{i+1},\dots,x_{n}].$
Homomorphisms between LAnKes are defined in the usual way. Following [5, Definition 2.1], the free LAnKe on a set $X$ is a LAnKe $\mathcal{L}$ together with a map $i:X\to\mathcal{L}$ such that if $f:X\to\mathcal{L^{\prime}}$ is a map, where $\mathcal{L^{\prime}}$ is a LAnKe, then there is a unique LAnKe homomorphism $F:\mathcal{L}\to\mathcal{L^{\prime}}$ such that $f=F\circ i$. By a standard argument, free LAnKes on $X$ are isomorphic. It is clear that the free LAnKe for $n=2$ is the free Lie algebra.
In [5], Friedmann, Hanlon, Stanley and Wachs initiated the study of the action of the symmetric group $\mathfrak{S}_{m}$ on the multilinear component of the free LAnKe. To be precise, the multilinear component $\text{Lie}_{n}(m)$ of the free LAnKe on $[m]:=\{1,\dots,m\}$ is spanned by the bracketed words on $[m]$ in which each $i$ appears exactly once. It follows that each such bracketed word has the same number of brackets, say $k$, and $m=(n-1)k+1$. Consider the action of $\mathfrak{S}_{m}$ on $\text{Lie}_{n}(m)$ given by replacing $i$ by $\sigma(i)$ in each bracketed word. Let us denote the corresponding representation of $\mathfrak{S}_{m}$ by $\rho_{n,k}$.
For a partition $\lambda$ of $m$, let $S^{\lambda}$ be the corresponding Specht module of the symmetric group $\mathfrak{S}_{m}$. As $\lambda$ ranges over all partitions of $m$, the modules $S^{\lambda}$ form a complete set of inequivalent representations of $\mathfrak{S}_{m}$. It was shown by Friedmann, Hanlon, Stanley and Wachs in [5, Theorem 1.3] that the representation $\rho_{n,2}$ of $\mathfrak{S}_{2n-1}$ is isomorphic to the Specht module $S^{(2^{n-1},1)}$ if $n\geq 2$. The following result is stated without proof in [5] and [4].
Theorem 1.2 ([4, Theorem 3],[5, p.4]).
The representation $\rho_{n,3}$ of $\mathfrak{S}_{3n-2}$ is isomorphic to the direct sum $S^{(3^{n-2},2,1^{2})}\oplus S^{(3^{n-1},1)}$ for every $n\geq 2$.
The decomposition of $\rho_{n,k}$ into irreducibles remains an open problem for $k\geq 4$.
The purpose of this note is to prove Theorem 1.2. We approach the problem within the framework of representations of the general linear group $G=GL_{N}(\mathbb{K})$. Using ideas from [10], which in turn were based on ideas of Brauner, Friedmann, Hanlon, Stanley and Wachs in [2, 5, 6], we define and study a particular $G$-equivariant map $M\oplus M\to M$, where $M=\Lambda^{n}V\otimes\Lambda^{n-1}V\otimes\Lambda^{n-1}V$ is the tensor product of the indicated exterior powers of the natural $G$-module of column vectors and $N\geq 3n-2$. This map has the property that when we apply the Schur functor, we obtain a presentation of the multilinear component $\text{Lie}_{n}(m)$ of the free LAnKe, where $m=3n-2$. We analyze the effect of the map on irreducible summands of $M$ using combinatorics of semistandard tableaux.
In Section 2 we establish notation and gather some recollections to be used in the sequel. In Section 3 we prove the main result for $G$ (see Theorem 3.3). In Section 4, we consider the setting of the symmetric group and show that Theorem 1.2 follows from Theorem 3.3.
2. Preliminaries
2.1. Notation
Let $G=GL_{N}(\mathbb{K})$. We denote by $V=\mathbb{K}^{N}$ be the natural $G$-module consisting of column vectors and by $D=\sum_{i\geq 0}D_{i}$ the divided power algebra of $V$, [12, Section 1.1]. We recall that $D$ is defined as the graded dual of the symmetric algebra of $V^{*}$, where $V^{*}$ is the dual of $V$. Since the characteristic of $\mathbb{K}$ is zero, $D$ is naturally isomorphic to the symmetric algebra $SV$ of $V$. However, the computations to be made in Section 3 seem less involved if one deals with Weyl modules in place of Schur modules. For this reason, we work with the divided power algebra and Weyl modules.
For a positive integer $r$, let $\Lambda(N,r)$ be the set of sequences $\alpha=(\alpha_{1},\dots,\alpha_{N})$ of length $N$ of nonnegative integers that $\alpha_{1}+\cdots+\alpha_{N}=r$ and let $\Lambda^{+}(N,r)$ be the subset of $\Lambda(N,r)$ consisting of partitions, that is sequences $\mu=(\mu_{1},\dots,\mu_{N})$ such that $\mu_{1}\geq\mu_{2}\geq\dots\geq\mu_{N}$. For $\alpha=(\alpha_{1},\dots,\alpha_{N})\in\Lambda(N,r)$, let $D(\alpha)=D(\alpha_{1},\dots,\alpha_{N})$ be the tensor product $D_{\alpha_{1}}\otimes\dots\otimes D_{\alpha_{N}}$ over $\mathbb{K}$. For $\mu\in\Lambda^{+}(N,r)$, we denote by $K_{\mu}$ the corresponding Weyl module for $G$ and by ${\pi}_{\mu}:D(\mu)\to K_{\mu}$ the natural projection. Since the characteristic of $\mathbb{K}$ is zero, the $G$-module $K_{\mu}$ is irreducible for every $\mu\in\Lambda^{+}(N,r)$. See [12, Section 2.2].
2.2. Tableaux and semistandard basis
Let us recall an important combinatorial property of $K_{\mu}$. We will restrict our discussion to partitions that have at most three parts.
We fix the order $e_{1}<e_{2}<...<e_{N}$ on the natural basis $\{e_{1},...,e_{N}\}$ of $V$. In the sequel we will denote $e_{i}$ by its subscript $i$. If $\mu=(\mu_{1},...,\mu_{N})\in\Lambda^{+}(N,r)$, a tableau of shape $\mu$ is a filling of the diagram of $\mu$ with entries from $\{1,...,N\}$. A tableau is called row semistandard if the entries are weakly increasing across the rows from left to right. A row semistandard tableau is called semistandard if the entries are strictly increasing down each column. We denote the set of semistandard tableaux of shape $\mu$ by $\mathrm{SST}(\mu)$. The weight of a tableau $S$ is the tuple $\alpha=(\alpha_{1},...,\alpha_{N})$, where $\alpha_{i}$ is the number of appearances of the entry $i$ in $S$. The set consisting of the semistandard (respectively, row semistandard) tableaux of shape $\mu$ and weight $\alpha$ will be denoted by $\mathrm{SST}_{\alpha}(\mu)$ (respectively, $\mathrm{RSST}_{\alpha}(\mu)$). For example, the following tableau of shape $\mu=(5,4,2)$
[…]
is semistandard and has weight $\alpha=(4,3,1,3)$. We will use ’exponential’ notation for row semistandard tableaux. If $S$ is a row semistandard tableau of shape $\mu$,
$S=\begin{matrix}[l]1^{(a_{1})}\cdots N^{(a_{N})}\\ 1^{(b_{1})}\cdots N^{(b_{N})}\\ 1^{(c_{1})}\cdots N^{(c_{N})},\end{matrix}$
where $a_{i},b_{i},c_{i}$ are nonnegative integers, let $e^{S}\in D(\mu)$ be the element
$e^{S}=1^{(a_{1})}\cdots N^{(a_{N})}\otimes 1^{(b_{1})}\cdots N^{(b_{N})}\otimes 1^{(c_{1})}\cdots N^{(c_{N})}.$
A classical result here is that there is a bijection between $\mathrm{SST}(\mu)$, and a basis of the $\mathbb{K}$-vector space $K_{\mu}$ given by $S\mapsto\pi_{\mu}(e^{S}).$ See [12, (2.1.15) Proposition]. We refer to the elements of this basis of $K_{\mu}$ as semistandard basis elements.
2.3. Projections
Since the characteristic of $\mathbb{K}$ is zero, every finite dimensional polynomial representation $M$ of $G$ is a direct sum of Weyl modules. The multiplicity of $K_{\mu}$, where $\mu\in\Lambda^{+}(N,r)$, as a summand of $M$ is equal to the dimension of the vector space $\operatorname{Hom}_{G}(M,K_{\mu})$. When $M=D(\alpha)$, where $\alpha\in\Lambda(N,r)$, the dimension of $\operatorname{Hom}_{G}(D(\alpha),K_{\mu})$, which is known as a Kostka number, is equal to the cardinality of the set $\mathrm{SST}_{\alpha}(\mu)$. In the sequel, we will need to identify different copies of $K_{\mu}$ in $D(\alpha)$. To this end we need an explicit basis of $\operatorname{Hom}_{G}(D(\alpha),K_{\mu})$ which we describe next.
Suppose $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\Lambda^{+}(3,r)$, $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})\in\Lambda(3,r)$ and $S$ is a row semistandard tableau of shape $\mu$ and weight $\alpha$
$S=\begin{matrix}[l]1^{(a_{11})}2^{(a_{12})}3^{(a_{13})}\\ 1^{(a_{21})}2^{(a_{22})}3^{(a_{23})}\\ 1^{(a_{31})}2^{(a_{32})}3^{(a_{33})}.\end{matrix}$
This means that for the matrix $A=(a_{ij})$ the row sums are given by $\mu$ and the column sums by $\alpha$. We refer to $A$ as the matrix of the row semistandard $S$. We define a map of $G$-modules $\phi_{S}:D(\alpha)\to D{(\mu)}$ as the following composition
(2.1) $D(\alpha_{1},\alpha_{2},\alpha_{3})\xrightarrow{\Delta_{1}\otimes\Delta_{2}\otimes\Delta_{3}}$ $D(a_{11},a_{21},a_{31})\otimes D(a_{12},a_{22},a_{32})\otimes D(a_{13},a_{23},a_{33})$
$\simeq$ $D(a_{11},a_{12},a_{13})\otimes D(a_{21},a_{22},a_{23})\otimes D(a_{31},a_{32},a_{33})$
$\xrightarrow{m_{1}\otimes m_{2}\otimes m_{3}}$ $D(\mu_{1},\mu_{2},\mu_{3}),$
where $\Delta_{i}:D(a_{1i}+a_{2i}+a_{3i})\to D(a_{1i},a_{2i},a_{3i})$ is the indicated component of threefold comultiplication of the Hopf algebra $D$, the isomorphism permutes tensor factors, and $m_{i}$: $D(a_{i1},a_{i2},a_{i3})\to D(a_{i1}+a_{i2}+a_{i3})$ is the indicated component of threefold multiplication in the algebra $D$.
In a similar way we have a map for exterior powers in place of divided powers. Define a map of $G$-modules $\psi_{S}:\Lambda(\alpha)\to\Lambda{(\mu)}$ as the following composition
(2.2) $\Lambda(\alpha_{1},\alpha_{2},\alpha_{3})\xrightarrow{\Delta_{1}\otimes\Delta_{2}\otimes\Delta_{3}}$ $\Lambda(a_{11},a_{21},a_{31})\otimes\Lambda(a_{12},a_{22},a_{32})\otimes\Lambda(a_{13},a_{23},a_{33})$
$\simeq$ $\Lambda(a_{11},a_{12},a_{13})\otimes\Lambda(a_{21},a_{22},a_{23})\otimes\Lambda(a_{31},a_{32},a_{33})$
$\xrightarrow{m_{1}\otimes m_{2}\otimes m_{3}}$ $\Lambda(\mu_{1},\mu_{2},\mu_{3}),$
where $\Delta_{i}$ is the indicated component of threefold comultiplication of the exterior algebra $\Lambda$, the isomorphism permutes tensor factors, and $m_{i}$ is the indicated component of threefold multiplication in the exterior algebra $\Lambda$.
Now we define $\pi_{S}:D(\alpha)\to K_{\mu}$ as the composition
$\pi_{S}:D(\alpha)\xrightarrow{\phi_{S}}D(\mu)\xrightarrow{\pi_{\mu}}K_{\mu}.$
We note that if the matrix $A$ is diagonal, then $\alpha=\mu$ and $\pi_{S}=\pi_{\mu}$. From [1, Section 2, equ. (11)] we know that a basis of the vector space $\operatorname{Hom}_{G}(D(\alpha),K_{\mu})$ is the set
(2.3) $\{\pi_{S}:S\in\mathrm{SST}_{\alpha}(\mu)\}.$
Remark 2.1.
It is well known that for every $\alpha\in\Lambda(N,r)$, the $G$-module $D(\alpha)$ is cyclic and a generator is the element $e^{\alpha}:=1^{(\alpha_{1})}\otimes 2^{(\alpha_{2})}\otimes\cdots\otimes N^{(\alpha_{N})}$. Hence $\pi_{S}$ is determined by the image $\pi_{S}(e^{\alpha})$ of $e^{\alpha}$.
3. Main result for Weyl modules
Throughout this section, $\lambda$ will be the partition $\lambda=(n,n-1,n-1)$.
3.1. The maps $\gamma$ and $\phi$
Definition 3.1.
Define $\gamma,\phi\in\operatorname{Hom}_{G}(D(\lambda),D(\lambda))$ by
$\gamma$ $=\phi_{S(1)}+(-1)^{n}\phi_{S(2)}$
$\phi$ $=\phi_{S(1)}+(-1)^{n}\phi_{S(3)}+(-1)^{n}\phi_{S(4)},$
where the tableaux $S(i)\in\mathrm{RSST}_{\lambda}(\lambda)$ are the following
$S(1)=\begin{matrix}[l]1^{(n)}\\ 2^{(n-1)}\\ 3^{(n-1)}\end{matrix},\;S(2)=\begin{matrix}[l]12^{(n-1)}\\ 1^{(n-1)}\\ 3^{(n-1)}\end{matrix},\;S(3)=\begin{matrix}[l]1^{(n)}\\ 3^{(n-1)}\\ 2^{(n-1)}\end{matrix},\;S(4)=\begin{matrix}[l]23^{(n-1)}\\ 12^{(n-2)}\\ 1^{(n-1)}\end{matrix}.$
Remark 3.2.
For short, let $\epsilon=(-1)^{n}$. We note that
(3.1) $\gamma(e^{\lambda})=e^{\lambda}+\epsilon 12^{(n-1)}\otimes 1^{(n-1)}\otimes 3^{(n-1)},$
(3.2) $\phi(e^{\lambda})=e^{\lambda}+\epsilon 1^{(n)}\otimes 3^{(n-1)}\otimes 2^{(n-1)}+\epsilon 23^{(n-1)}\otimes 12^{(n-2)}\otimes 1^{(n-1)}.$
The motivation for considering such maps will become clear in sections 4.1 and 4.2. Roughly speaking, the images of these maps correspond to relations of $\text{Lie}_{n}(m)$ that are consequences of the generalized Jacobi identity.
We intend to prove the following result.
Theorem 3.3.
Let $N\geq 3n-2$. Then, the cokernel of the map
$\gamma+\phi:D(\lambda)\oplus D(\lambda)\to D(\lambda)$
is isomorphic to $K_{\lambda}\oplus K_{(n+1,n-1,n-2)}$ as $G$-modules.
3.2. The actions of the maps $\gamma$ and $\phi$
The next four lemmas analyze the action of the map $\gamma+\phi$ on the irreducible summands of $D(\lambda)$. The first two lemmas concern the summands $K_{\lambda}$ and $K_{(n+1,n-1,n-2)}$. The last two lemmas show that the other possible summands have a combinatorial property that will be utilized is Section 3.3 to prove that their multiplicities in $\operatorname{Coker}(\gamma+\phi)$ are in fact zero.
Lemma 3.4.
The multiplicity of $K_{\lambda}$ in $\operatorname{Coker}{(\gamma+\phi)}$ is equal to one.
Proof.
Let $T_{0}=\pi_{\lambda}(e^{\lambda})$. Using (3.1) we have
$\pi_{\lambda}\circ\gamma(e^{\lambda})=T_{0}+(-1)^{n}\pi_{\lambda}(12^{(n-1)}\otimes 1^{(n-1)}\otimes 3^{(n-1)}).$
Applying [11, Lemma 4.2(2)] to raise the 1’s from the second row to the first, we have $\pi_{\lambda}(12^{(n-1)}\otimes 1^{(n-1)}\otimes 3^{(n-1)})=(-1)^{n-1}T_{0}$. By substituting we obtain $\pi_{\lambda}\circ\gamma(e^{\mu})=T_{0}+(-1)^{2n-1}T_{0}=0.$ A similar computation using (3.2) yields
$\pi_{\lambda}\circ\phi(e^{\lambda})$ $=T_{0}+(-1)^{n}\pi_{\lambda}(1^{(n)}\otimes 3^{(n-1)}\otimes 2^{(n-1)})$
$\;\;\;\;+(-1)^{n}\pi_{\lambda}(23^{(n-1)}\otimes 12^{(n-1)}\otimes 1^{(n-1)})$
$=T_{0}+(-1)^{2n-1}T_{0}+0$
$=0,$
where in the second equality we used [11, Lemma 4.2(2)] and [11, Lemma 4.2(1)]. Thus $K_{\lambda}$ is not a summand of the image $\operatorname{Im}(\gamma+\phi)$. Since the multiplicity of $K_{\lambda}$ in the codomain $D({\lambda})$ of the map $\gamma+\phi$ is equal to 1, the result follows. ∎
Lemma 3.5.
The multiplicity of $K_{(n+1,n-1,n-2)}$ in $\operatorname{Coker}{(\gamma+\phi)}$ is equal to one.
Proof.
Let $\mu=(n+1,n-1,n-2)$. There are exactly two semistandard tableaux of shape $\mu$ and weight $\lambda$,
$S_{0}=\begin{matrix}[l]1^{(n)}3\\ 2^{(n-1)}\\ 3^{(n-2)}\end{matrix},\;S_{1}=\begin{matrix}[l]1^{(n)}2\\ 2^{(n-2)}3\\ 3^{(n-2)}\end{matrix}.$
According to Section 2.3, the corresponding projections are
$\pi_{S_{0}}:D(\lambda)\to K_{\mu},e^{\lambda}\mapsto\pi_{\mu}(1^{(n)}3\otimes 2^{(n-1)}\otimes 3^{(n-2)}),$
$\pi_{S_{1}}:D(\lambda)\to K_{\mu},e^{\lambda}\mapsto\pi_{\mu}(1^{(n)}2\otimes 2^{(n-2)}3\otimes 3^{(n-2)}).$
We compute the maps $\pi_{S_{i}}\circ\gamma,\pi_{S_{i}}\circ\phi\in\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$. For short, let $T_{i}=\pi_{S_{i}}(e^{\lambda})$, $i=0,1$. Using (3.1) in the first of the next equalities and [11, Lemma 4.2(2)] in the second, we obtain
$\pi_{S_{0}}\circ\gamma(e^{\lambda})$ $=T_{0}+(-1)^{n}\pi_{\mu}(12^{(n-1)}3\otimes 1^{(n-1)}\otimes 3^{(n-2)})$
$=T_{0}+(-1)^{2n-1}\big{(}\pi_{\mu}(1^{(n)}3\otimes 2^{(n-1)}\otimes 3^{(n-2)})$
$\;\;\;\;+\pi_{\mu}(1^{(n)}2\otimes 2^{(n-2)}3\otimes 3^{(n-2)})\big{)}$
$=T_{0}-T_{0}-T_{1}$
$=-T_{1}.$
In a similar manner we have
$\pi_{S_{1}}\circ\gamma(e^{\lambda})$ $=T_{1}+(-1)^{n}\tbinom{2}{1}\pi_{\mu}(1^{(2)}2^{(n-1)}\otimes 1^{(n-2)}3\otimes 3^{(n-2)})$
$=T_{1}+(-1)^{2n-2}\tbinom{2}{1}T_{1}$
$=3T_{1}.$
Since $e^{\lambda}$ generates the $G$-module $D(\lambda)$, we conclude from the above that the restrictions of the maps $\pi_{S_{0}},\pi_{S_{1}}\in\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$ to the image $\operatorname{Im}(\gamma)\subseteq D(\lambda)$ are linearly dependent. We show below that the restrictions of $\pi_{S_{0}},\pi_{S_{1}}$ to the image $\operatorname{Im}(\phi)\subseteq D(\lambda)$ are both zero. This implies that the restrictions of $\pi_{S_{0}},\pi_{S_{1}}$ to $\operatorname{Im}(\gamma+\phi)$ are linearly dependent. By the above computation, these restrictions are nonzero since $T_{1}\neq 0$. Thus the multiplicity of $K_{\mu}$ in $\operatorname{Im}(\gamma+\phi)$ is equal to 1. Finally, the multiplicity of $K_{\mu}$ in $\operatorname{Coker}(\gamma+\phi)$ is equal to $2-1=1$. It remains to be shown that the restrictions of $\pi_{S_{0}},\pi_{S_{1}}$ to the image $\operatorname{Im}(\phi)\subseteq D(\lambda)$ are both zero. To this end, we first show the identity
(3.3) $\pi_{\mu}(123^{(n-1)}\otimes 12^{(n-2)}\otimes 1^{(n-2)})=(-1)^{n-1}(T_{0}+T_{1}).$
Indeed, applying [11, Lemma 4.2(2)] several times starting from rows 2 and 3, we have
$\pi_{\mu}(123^{(n-1)}\otimes 12^{(n-2)}\otimes 1^{(n-2)})$ $=(-1)^{n-2}\pi_{\mu}(123^{(n-1)}\otimes 1^{(n-1)}\otimes 2^{(n-2)})$
$=-\big{(}\pi_{\mu}(1^{(n)}2\otimes 3^{(n-1)}\otimes 2^{(n-2)})$
$\;\;\;\;+\pi_{\mu}(1^{(n)}3\otimes 23^{(n-2)}\otimes 2^{(n-2)})\big{)}$
$=(-1)^{n-1}(T_{0}+T_{1}).$
Now using (3.2) we obtain
$\pi_{S_{0}}\circ\phi(e^{\lambda})$ $=T_{0}+(-1)^{n}\pi_{\mu}(1^{(n)}2\otimes 3^{(n-1)}\otimes 2^{(n-2)})$
$\;\;\;\;+(-1)^{n}\pi_{\mu}(123^{(n-1)}\otimes 12^{(n-2)}\otimes 1^{(n-2)})$
$=T_{0}+(-1)^{n+n-2}T_{1}+(-1)^{n+n-1}(T_{0}+T_{1})$
$=0,$
where in the second equality we used once again [11, Lemma 4.2(2)] and (3.3). By a similar computation, we have
$\pi_{S_{1}}\circ\phi(e^{\lambda})$ $=T_{1}+(-1)^{n}\pi_{\mu}(1^{(n)}3\otimes 23^{(n-2)}\otimes 2^{(n-2)})$
$\;\;\;\;+(-1)^{n}\Big{(}\pi_{\mu}(123^{(n-1)}\otimes 12^{(n-2)}\otimes 1^{(n-2)})$
$\;\;\;\;+\tbinom{2}{1}^{2}\pi_{\mu}(2^{(2)}3^{(n-1)}\otimes 1^{(2)}2^{(n-3)}\otimes 1^{(n-2)})\Big{)}$
$=T_{1}+(-1)^{n+n-2}T_{0}+(-1)^{n+n-1}(T_{0}+T_{1}+\tbinom{2}{1}^{2}0)$
$=0,$
where we used $\pi_{\mu}(2^{(2)}3^{(n-1)}\otimes 1^{(2)}2^{(n-3)}\otimes 1^{(n-2)})=0$ according to [11, Lemma 4.2(1)] applied to rows 2 and 3. ∎
The next two lemmas concern irreducible summands of the tensor product $K_{(n,n-1)}\otimes D_{n-1}$. We need some notation. Let $\mathrm{Par}(K_{(n,n-1)}\otimes D_{n-1})$ be the subset of $\Lambda(N,3n-2)$ consisting of the partitions $\mu$ such that $K_{\mu}$ is a summand of $K_{(n,n-1)}\otimes D_{n-1}$. By Pieri’s rule, for example see [12, Corollary 2.3.5], we have that $\mathrm{Par}(K_{(n,n-1)}\otimes D_{n-1})$ consists of those partitions $\mu$ that are of the form
(3.4) $\mu=(n+c_{1},n-1+c_{2},c_{3}),$
for some nonnegative integers $c_{1},c_{2},c_{3}$ satisfying $c_{1}+c_{2}+c_{3}=n-1$ and $c_{2}\in\{0,1\}$.
Let us fix a partition $\mu=(n+c_{1},n-1+c_{2},n-1-c_{1}-c_{2})$ as in (3.4). One easily verifies that for each $i=0,\dots,c_{1}$, there is a unique semistandard tableau $S_{i}\in\mathrm{SST}_{\lambda}(\mu)$ such that the number of 2’s in the first row is equal to $i$, and moreover $\mathrm{SST}_{\lambda}(\mu)=\{S_{0},S_{1},\dots,S_{c_{1}}\}.$ We have
(3.5) $S_{i}=\begin{matrix}[l]1^{(n)}2^{(i)}3^{(c_{1}-i)}\\ 2^{(n-1-i)}3^{(c_{2}+i)}\\ 3^{(n-1-c_{1}-c_{2})}.\end{matrix}$
The next lemma concerns the action of $\gamma$ on $K_{\mu}$.
Lemma 3.6.
Let $\mu\in\mathrm{Par}(K_{(n,n-1)}\otimes D_{n-1})$. Then for every $S_{i}\in\mathrm{SST}_{\lambda}(\mu)$, the coefficient of $T_{0}:=\pi_{\mu}(e^{S_{0}})\in K_{\mu}$ in the expression of $\pi_{S_{i}}\circ\gamma(e^{\lambda})\in K_{\mu}$ as a linear combination of semistandard basis elements is equal to 0.
Proof.
Let $\mu=(n+c_{1},n-1+c_{2},n-1-c_{1}-c_{2})$ according to (3.4). For short, let $c=c_{1}+c_{2}$. Also, let $T_{i}=\pi_{\mu}(e^{S_{i}})$, $i=0,\dots,c_{1}$. Using (3.1) and [11, Lemma 4.2(2)] for rows 1 and 2 we compute
$\pi_{S_{i}}\circ\gamma(e^{\lambda})$ $=T_{i}+(-1)^{n}\tbinom{i+1}{1}\pi_{\mu}(1^{(i+1)}2^{(n-1)}3^{(c_{1}-i)}\otimes 1^{(n-1-i)}3^{(c_{2}+i)}\otimes 3^{(n-1-c)})$
$=T_{i}+(-1)^{i+1}\tbinom{i+1}{1}\big{(}T_{i}+\tbinom{c_{2}+i+1}{1}T_{i+1}+\cdots+\tbinom{c_{1}+c_{2}}{c_{1}-i}T_{c_{1}}\big{)}.$
If $i>0$, the coefficient of $T_{0}$ in the right hand side is equal to 0. If $i=0$, the coefficient of $T_{0}$ in the right hand side is equal to $1+(-1)=0$. ∎
We keep the previous notation. The next lemma concerns the action of $\phi$ on $K_{\mu}$.
Lemma 3.7.
Let $\mu\in\mathrm{Par}(K_{(n,n-1)}\otimes D_{n})$ such that $\mu\neq\lambda,(n+1,n-1,n-2)$. Then there is an $S\in\mathrm{SST}_{\lambda}(\mu)$ such that the coefficient of $T_{0}:=\pi_{\mu}(e^{S_{0}})\in K_{\mu}$ in the expression of $\pi_{S}\circ\phi(e^{\lambda})\in K_{\mu}$ as a linear combination of semistandard basis elements is nonzero.
Proof.
We know that $\mu$ is of the form $\mu=(n+c_{1},n-1+c_{2},n-1-c_{1}-c_{2})$ by (3.4). Let $S=S_{0}$, which is defined in (3.5). For short, let $c=c_{1}+c_{2}$. We compute using (3.2) and [11, Lemma 4.2(2)]
$\pi_{S}\circ\phi(e^{\lambda})$ $=T_{0}+(-1)^{n}\pi_{\mu}(1^{(n)}2^{(c_{1})}\otimes 2^{(c_{2})}3^{(n-1)}+2^{(n-1-c)})$
$\;\;\;\;+(-1)^{n}\tbinom{1+c_{2}}{1}\pi_{\mu}(1^{(c_{1})}23^{(n-1)}\otimes 1^{(1+c_{2})}2^{(n-2)}\otimes 1^{(n-1-c)})$
$=T_{0}+(-1)^{c+1}T_{c_{1}}$
$\;\;\;\;+(-1)^{c+1}\tbinom{1+c_{2}}{1}\pi_{\mu}(1^{(c_{1})}23^{(n-1)}\otimes 1^{(n-c_{1})}2^{(c-1)}\otimes 2^{(n-1-c))}.$
In order to continue the computation we distinguish two cases. Suppose $c_{1}\geq 1$. Then by raising the 1’s from row 2 according to [11, Lemma 4.2(2)] and raising the 2’s from row 3 we obtain
$\pi_{S}\circ\phi(e^{\lambda})$ $=T_{0}+(-1)^{c+1}T_{c_{1}}$
$\;\;\;\;+(-1)^{n+c_{2}+1}\tbinom{1+c_{2}}{1}\Big{(}\pi_{\mu}(1^{(n)}23^{(c_{1}-1)}\otimes 2^{(c-1)}3^{(n-c_{1})}\otimes 2^{(n-1-c)})$
$\;\;\;\;+\tbinom{c}{1}\pi_{\mu}(1^{(n)}3^{(c_{1})}\otimes 2^{(c)}3^{(n-c_{1}-1)}\otimes 2^{(n-1-c)})\Big{)}$
$=T_{0}+(-1)^{c+1}T_{c_{1}}+(-1)^{c_{1}}\tbinom{1+c_{2}}{1}(T_{1}+\tbinom{c}{1}T_{0}).$
The coefficient of $T_{0}$ in the right hand side is equal to $1+(-1)^{c_{1}}\tbinom{1+c_{2}}{1}\tbinom{c}{1}.$ This is clearly nonzero if $c_{2}=1$. If $c_{2}=0$, then the coefficient is equal to $1+(-1)^{c_{1}}c_{1}$. However, for $c_{2}=0$ we have $c_{1}\neq 1$ because $\mu\neq(n+1,n-1,n-2)$ by hypothesis. Thus we see that the coefficient is nonzero when $c_{2}=0$. (Remember that $c_{2}=0,1$.) Suppose $c_{1}=0$. Then by raising the 1’s from row 2 and the 2’s from row 3 as above, we obtain
$\pi_{S}\circ\phi(e^{\lambda})$ $=T_{0}+(-1)^{c_{2}+1}T_{0}$
$\;\;\;\;+(-1)^{n+c_{2}+1}\tbinom{1+c_{2}}{1}\Big{(}\tbinom{c_{2}}{1}\pi_{\mu}(1^{(n)}\otimes 2^{(c_{2})}3^{(n-1)}\otimes 2^{(n-1-c_{2})})\Big{)}$
$=(1+(-1)^{c_{2}+1}+\tbinom{1+c_{2}}{1}\tbinom{c_{2}}{1})T_{0}.$
If $c_{2}=1$, the coefficient of $T_{0}$ is nonzero. We have $c_{2}\neq 0$ because $\mu\neq\lambda=(n,n-1,n-1)$ by hypothesis. ∎
3.3. The functor $\Omega$
With the previous four lemmas we may prove Theorem 3.3.
Let us recall that there is an algebra involution on the ring of symmetric functions that sends the Schur function $s_{\mu}$ to $s_{\mu^{\prime}}$ for every partition $\mu$ [8, 6.2]. In terms of representations, we recall from [1, p. 189] that, since the characteristic of $\mathbb{K}$ is zero, there is an involutory natural equivalence $\Omega$ from the category of polynomial representations of $G$ of degree $r$ to itself that has the following properties.
  1. (1)
    $\Omega(D(\alpha))=\Lambda(\alpha)$ for all $\alpha=(\alpha_{1},\dots,a_{N})\in\Lambda(N,r)$, where $\Lambda(\alpha)=\Lambda^{\alpha_{1}}\otimes\cdots\otimes\Lambda^{\alpha_{N}}$ is the indicated tensor product of exterior powers of the natural $G$-module $V$.
  2. (2)
    $\Omega(K_{\mu})=L_{\mu}$ for every $\mu\in\Lambda^{+}(N,r)$, where $L_{\mu}$ is the Schur module corresponding to $\mu$ defined in [12, Section 2.1].
  3. (3)
    $\Omega$ preserves the comultiplication and multiplication maps of the Hopf algebras $D$ and $\Lambda$.
  4. (4)
    $\Omega(\tau_{i,D})=(-1)^{\alpha_{i}a_{i+1}}\tau_{i,\Lambda}$ for all $\alpha=(\alpha_{1},\dots,a_{N})\in\Lambda(N,r)$, where $\tau_{i,D}:D(\alpha)\to D(\alpha)$ (respectively, $\tau_{i,\Lambda}:\Lambda(\alpha)\to\Lambda(\alpha)$ is the map that interchanges the factors $D_{a_{i}}$ and $D_{a_{i+1}}$ (respectively, $\Lambda^{a_{i}}$ and $\Lambda^{a_{i+1}}$) and is the identity on the rest.
3.4. Proof of Theorem 3.3
Proof.
From Lemma 3.4 and Lemma 3.5 we know that the multiplicity of each of $K_{\lambda}$ and $K_{(n+1,n-1,n-2)}$ in the cokernel of the map $\gamma+\phi:D(\lambda)\oplus D(\lambda)\to D(\lambda)$ is equal to 1. Applying the functor $\Omega$ to [10, Corollary 5.4] for $k=n-1$, it follows that the cokernel of the map $\gamma:D(\lambda)\to D(\lambda)$ is
(3.6) $\operatorname{Coker}(\gamma)\simeq K_{(n,n-1)}\otimes D_{n-1}.$
(This also follows from [5, Theorem 1.3] by applying first the inverse Schur functor and then the duality functor $\Omega$.) Let us fix $\mu\in\mathrm{Par}(K_{(n,n-1)}\otimes D_{n-1})$ such that $\mu\neq\lambda$ and $\mu\neq(n+1,n-1,n-2).$ We intend to show that the multiplicity of $K_{\mu}$ in $D(\lambda)$ is equal to the multiplicity of $K_{\mu}$ in $\operatorname{Im}(\gamma)+\operatorname{Im}(\phi)$, or equivalently, that the vector spaces $\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$ and $\operatorname{Hom}_{G}(\operatorname{Im}(\gamma)+\operatorname{Im}(\phi),K_{\mu})$ have equal dimensions. We need to recall some notation. Let $\mathrm{SST_{\lambda}}(\mu)=\{S_{0},S_{1},\cdots,S_{q}\},$ where $S_{i}$ is given by (3.5). For each $S_{i}$ we have the corresponding projection $\pi_{S_{i}}:D(\lambda)\to K_{\mu},e^{\lambda}\mapsto\pi_{\mu}(e^{S_{i}})$, and we know that the $\pi_{S_{0}},\dots,\pi_{S_{q}}$ form a basis of $\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$. The exact sequence $0\to\operatorname{Im}(\gamma)\to D(\lambda)\to\operatorname{Coker}(\gamma)\to 0$ yields the exact sequence
(3.7) $0\to\operatorname{Hom}_{G}(\operatorname{Coker}(\gamma),K_{\mu})\to\operatorname{Hom}_{G}(D(\lambda),K_{\mu})\to\operatorname{Hom}_{G}(\operatorname{Im}(\gamma),K_{\mu})\to 0$
From (3.6) and Pieri’s rule, it follows that $\dim\operatorname{Hom}_{G}(\operatorname{Coker}(\gamma),K_{\mu})=1$ and hence from (3.7) we conclude that $\dim\operatorname{Hom}_{G}(\operatorname{Im}(\gamma),K_{\mu})=q$. We have that $\operatorname{Hom}_{G}(\operatorname{Im}(\gamma),K_{\mu})$ is generated by the restrictions $\pi_{S_{i}}\big{|}_{\operatorname{Im}{(\gamma)}}$, $i=0,\dots,q$, of the maps $\pi_{S_{i}}$ to $\operatorname{Im}(\gamma).$ Consider the maps
$\pi_{S_{0}}\circ\gamma,\pi_{S_{1}}\circ\gamma,\cdots,\pi_{S_{q}}\circ\gamma\in\operatorname{Hom}_{G}(D(\lambda),K_{\mu}).$
It is clear that if $a_{0}\pi_{S_{0}}\circ\gamma+\cdots+a_{q}\pi_{S_{q}}\circ\gamma=0,$ where $a_{i}\in\mathbb{K}$, then $a_{0}\pi_{S_{0}}\big{|}_{\operatorname{Im}{(\gamma)}}+\cdots+a_{q}\pi_{S_{q}}\big{|}_{\operatorname{Im}{(\gamma)}}=0.$ Thus, for the subspace
$U=\mathrm{span}\{\pi_{S_{0}}\circ\gamma,\pi_{S_{1}}\circ\gamma,\cdots,\pi_{S_{q}}\circ\gamma\}$
of $\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$ we have
$\dim U\geq\dim\operatorname{Hom}_{G}(\operatorname{Im}(\gamma),K_{\mu})=q.$
We claim that the subspace
$\mathrm{span}\{\pi_{S_{0}}\circ\phi,\pi_{S_{0}}\circ\gamma,\pi_{S_{1}}\circ\gamma,\cdots,\pi_{S_{q}}\circ\gamma\}$
of $\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$ has dimension $q+1$. Indeed, to prove this it suffices to show that $\pi_{S_{0}}\circ\phi\notin U.$ Suppose $\pi_{S_{0}}\circ\phi=a_{0}\pi_{S_{0}}\circ\gamma+\cdots+a_{q}\pi_{S_{q}}\circ\gamma$, where $a_{i}\in\mathbb{K}$. Evaluating at $e^{\lambda}$ we have
$\pi_{S_{0}}\circ\phi(e^{\lambda})=a_{0}\pi_{S_{0}}\circ\gamma(e^{\lambda})+\cdots+a_{q}\pi_{S_{q}}\circ\gamma(e^{\lambda}).$
According to Lemma 3.7, the coefficient of the semistandard basis element $T_{0}=\pi_{\mu}(e^{S_{0}})$ in the left hand side is nonzero, and according to Lemma 3.6 the coefficient of $T_{0}$ in the right hand side is zero. Thus $\pi_{S_{0}}\circ\phi\notin U$ as desired. It follows from the above claim that there is a basis of $\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$ consisting of a subset of the elements
(3.8) $\pi_{S_{0}}\circ\phi,\pi_{S_{0}}\circ\gamma,\pi_{S_{1}}\circ\gamma,\cdots,\pi_{S_{q}}\circ\gamma.$
Since every map in (3.8) factors through a subspace of $\operatorname{Im}(\gamma)+\operatorname{Im}(\phi)$, we conclude that the multiplicity of $K_{\mu}$ in $\operatorname{Im}(\gamma)+\operatorname{Im}(\phi)\subseteq D(\lambda)$ is at least $\dim\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$. Thus this multiplicity is equal to $\dim\operatorname{Hom}_{G}(D(\lambda),K_{\mu})$. Consequently, the multiplicity of $K_{\mu}$ in $\operatorname{Coker}(\gamma+\phi)$ is equal to zero. ∎
4. LAnKes and Specht modules
Let $m=3n-2$. The purpose of this section is to show how Theorem 1.2 follows from Theorem 3.3. To this end, we intend (a) to apply to Theorem 3.3 the functor $\Omega$ and then the Schur functor, and (b) to identify the relations in the resulting quotient in order to show that it is isomorphic to $\text{Lie}_{n}(m)$ as an $\mathfrak{S}_{m}$-module. But first, let us make explicit a presentation of $\text{Lie}_{n}(m)$.
4.1. A presentation of $\text{Lie}_{n}(m)$
Consider a bracketed word in $\text{Lie}_{n}(m)$. The number of brackets is $k=3$. Hence up to skew commutativity of the bracket, there are two type of generators of $\text{Lie}_{n}(m)$ consisting of bracketed words
(G1) $[[[{x_{1}},\dots,{x_{n}}],{y_{1}},\dots,y_{n-1}],z_{1},\dots,z_{n-1}]$
(G2) $[[{x_{1}},\dots,{x_{n}}],[{y_{1}},\dots,y_{n}],z_{1},\dots,z_{n-2}],$
where in both cases $x_{i},y_{j},z_{k}\in[m]$ run over all choices such that in each of (G1) and (G2) every $a\in[m]$ appears exactly once. By skew commutativity of the bracket, we may assume for (G1) that $x_{1}<\dots<x_{n}$, $y_{1}<\dots<y_{n-1}$, $z_{1}<\dots<z_{n-1}$ and for (G2) that $x_{1}<\dots<x_{n}$, $y_{1}<\dots<y_{n}$, $z_{1}<\dots<z_{n-2}$ and $x_{1}<y_{1}$. The relations among these are consequences of the generalized Jacobi identity. This identity may be applied in exactly two ways to a generator of type (G1); we may exchange the $y_{1},\dots,y_{n-1}$ with $n-1$ of the $x_{1},\dots,x_{n}$ (keeping the $z_{k}$ fixed), or we may exchange the $z_{1},\dots,z_{n-1}$ with $n-1$ of the $[x_{1},\dots,x_{n}],y_{1},\dots,y_{n-1}$. Thus we have the relations (R1) and (R2) below. The generalized Jacobi identity may be applied in exactly one way to a generator of type (G2), by exchanging the $[y_{1},\dots,y_{n}],z_{1},\dots,z_{n-2}$ with $n-1$ of the $x_{1},\dots,x_{n}$. Thus we have the relation (R3) below.
(R1) (G1) $-\sum_{i=1}^{n}[[{x_{1}},\dots,{x_{i-1}},[x_{i},{y_{1}},\dots,y_{n-1}],x_{i+1},\dots,x_{n}],z_{1},\dots,z_{n-1}],$
(R2) (G1) $-[[[{x_{1}},\dots,{x_{n}}],{z_{1}},\dots,z_{n-1}],y_{1},\dots,y_{n-1}]$
$-\sum_{i=1}^{n-1}[[{x_{1}},\dots,{x_{n}}],{y_{1}},\dots,y_{i-1},[y_{i},z_{1},\dots,x_{n-1}],y_{i+1},\dots,y_{n-1}],$
(R3) (G2) $-\sum_{i=1}^{n}[x_{1},\dots,x_{i-1},[x_{i},[y_{1},\dots,y_{n}],z_{1},\dots,z_{n-2}],x_{i+1},\dots,x_{n}].$
Hence we have a presentation of $\text{Lie}_{n}(m)$ with generators (G1), (G2) and relations (R1), (R2), (R3).
Note that by skew commutativity of the bracket, every summand in the sum of (R2) is up to sign a generator of type (G2). Also, every summand in the sum of (R3) is up to sign a generator of type (G1). We substitute (R3) in (R2) to obtain
(R4) (G1) $-[[[{x_{1}},\dots,{x_{n}}],{z_{1}},\dots,z_{n-1}],y_{1},\dots,y_{n-1}]$
$-\sum_{i=1}^{n-1}(-1)^{i-1}\sum_{j=1}^{n}(-1)^{j}[[[y_{i},z_{1},\dots,z_{n-1}],x_{j},y_{1},\dots,\hat{y}_{i},\dots,y_{n-1}],$
$x_{1},\dots,\hat{x}_{j},\dots,x_{n}].$
Hence we have shown the following lemma.
Lemma 4.1.
[…]
4.2. A presentation associated to the map $\gamma+\phi$
For the remainder of this section, let $\lambda=(n,n-1,n-1)$ and $\mu=(n+1,n-1,n-2)$.
Applying the functor $\Omega$ that we discussed in Section 3.3 to the map $\gamma+\phi$ of Theorem 3.3, we obtain a map $G$-modules
(4.1) $\Lambda(\lambda)\oplus\Lambda(\lambda)\xrightarrow{\Omega(\gamma)+\Omega(\phi)}\Lambda(\lambda).$
From Definition 3.1, the definition of $\phi_{S}$ and $\psi_{S}$ in (2.1) and (2.2), and the description of the functor $\Omega$ (especially item 4 that involves sign changes), it is straightforward to verify that the maps $\Omega(\gamma)$ and $\Omega(\phi)$ are as follows,
$\Omega(\gamma)$ $=1_{\Lambda(\lambda)}-\psi_{S_{2}}$
$\Omega(\phi)$ $=1_{\Lambda(\lambda)}-\psi_{S_{3}}+\psi_{S_{4}}.$
We will need to compute these maps on basis elements of $\Lambda(\lambda)$. Recall that we have the basis $\{e_{1},\dots,e_{N}\}$ of the natural $G$-module $V$. So let
$w=x\otimes y\otimes z=x_{1}\cdots x_{n}\otimes y_{1}\cdots y_{n-1}\otimes z_{1}\cdots z_{n-1}\in\Lambda(\lambda),$
where $x_{i},y_{j},z_{j}\in\{e_{1},\dots,e_{N}\}$. For $i\in\{1,\dots,n\}$ we let
$x[i]=x_{1}\cdots\hat{x_{i}}\cdots x_{n}\in\Lambda^{n-1},$
where $\hat{x_{i}}$ means that $x_{i}$ is omitted. For $j\in\{1,\dots,n-1\}$ we define $y[j]\in\Lambda^{n-2}$ in a similar manner. It is easy to check using the definition of $\psi_{S}$ in (2.2) and the definition of the tableaux $S_{i}$ in Definition 3.8, that
(4.2) $\Omega(\gamma)(w)$ $=w-\sum_{i=1}^{n}(-1)^{i-1}x_{i}y\otimes x[i]\otimes z$
(4.3) $\Omega(\phi)(w)$ $=w-x\otimes z\otimes y+\sum_{i=1}^{n-1}(-1)^{i-1}\sum_{j=1}^{n}(-1)^{j-1}y_{i}z\otimes x_{j}y[i]\otimes x[j].$
Next we want to apply the Schur functor. Suppose $N\geq m$. Recall from [9], the Schur functor is a functor $f$ from the category of homogeneous polynomial representations of $G$ of degree $m$ to the category of left $\mathfrak{S}_{m}$-modules. For $M$ an object in the first category, $f(M)$ is the weight subspace $M_{\alpha}$ of $M$, where $\alpha=(1^{m},0^{N-m})$ and for $\theta:M\to N$ a morphism in the first category, $f(\theta)$ is the restriction $M_{\alpha}\to N_{\alpha}$ of $\theta$. Let us denote the conjugate of a partition $\nu$ by $\nu^{\prime}$. The Specht module corresponding to a partition $\nu$ of $m$ will be denoted by $S^{\nu}$ and the space of column tabloids corresponding to $\nu$ will be denoted by ${\tilde{M}}^{\nu}$. See [8, Chapter 7.4] for the later. It is well known that $f$ is an exact functor such that $f(L_{\nu})=S^{\nu^{\prime}}$ and $f(\Lambda^{\nu})=\tilde{M}^{\nu^{\prime}}$.
Applying the Schur functor to the map $\Lambda(\lambda)\oplus\Lambda(\lambda)\xrightarrow{\Omega(\gamma)+\Omega(\phi)}\Lambda(\lambda)$ we obtain a map of $\mathfrak{S}_{m}$-modules
(4.4) $\tilde{M}^{\lambda^{\prime}}\oplus\tilde{M}^{\lambda^{\prime}}\xrightarrow{f(\Omega(\gamma))+f(\Omega(\phi))}\tilde{M}^{\lambda^{\prime}}.$
If an element $x\otimes y\otimes z$ of $\Lambda(\lambda)$ is in $\tilde{M}^{\lambda^{\prime}}$, we denote by $x|y|z$ its image in the cokernel of (4.4).
Lemma 4.2.
The cokernel of the map (4.4) has a presentation with generators $x|y|z$ and relations
(4.5) $x|y|z=\sum_{i=1}^{n}(-1)^{i-1}x_{i}y|x[i]|z,$
(4.6) $x|y|z=x|z|y-\sum_{i=1}^{n-1}(-1)^{i-1}\sum_{j=1}^{n}(-1)^{j-1}y_{i}z|x_{j}y[i]|x[j],$
where $x\in\Lambda^{n}$ and $y,z\in\Lambda^{n-1}$ run over all elements of the form $x=x_{1}\cdots x_{n}$, $y=y_{1}\cdots y_{{n-1}}$, $z=z_{1}\cdots z_{n-1}$ such that $x_{1},\dots x_{n},y_{1},\dots,y_{n-1},z_{1},\dots,z_{n-1}$ is permutation of $e_{1},e_{2},\dots,e_{3n-2}.$
Proof.
This follows from the previous discussion and the equalities (4.2), (4.3). ∎
4.3. Proof of Theorem 1.2
Recall that we are assuming $N\geq m=3n-2$ and $n\geq 2$.
Consider the map of $\mathfrak{S}_{m}$-modules
$h:\tilde{M}^{\lambda^{\prime}}$ $\to\text{Lie}_{n}(m),$
$x_{1}\cdots x_{n}|y_{1}\cdots y_{n-1}|z_{1}\cdots z_{n-1}$ $\mapsto[[[x_{1},\dots,x_{n}],y_{1},\dots,y_{n-1}],z_{1},\dots,z_{n-1}].$
From Lemma 4.2 and Lemma 4.1 it follows that $h$ induces a map of $\mathfrak{S}_{m}$-modules
$\operatorname{Coker}(\tilde{M}^{\lambda^{\prime}}\oplus\tilde{M}^{\lambda^{\prime}}\xrightarrow{f(\Omega(\gamma))+f(\Omega(\phi))}\tilde{M}^{\lambda^{\prime}})\to\text{Lie}_{n}(m)$
which is an isomorphism, since $h$ carries (4.5) and (4.6) to (R1) and (R4) respectively. Hence, from Theorem 3.3 it follows that $\text{Lie}_{n}(m)\simeq S^{\lambda^{\prime}}\oplus S^{\mu^{\prime}}$ as $\mathfrak{S}_{m}$-modules.